sonteya:

The vertices of triangle JKL, shown below, are J(2, 1), K(4, 4), and L(6, 2). If triangle JKL is translated 2 units to the left and 4 units down, what are the new coordinates of Point L?

carmelle:

Rules: Left 2 → (x - 2) Down 4 → (y - 4) Now, let's find L. L (6, 2) → (6 - 2, 2 - 4) We solve by subtracting... 6 - 2 = 4 ( `x` , y ) 2 - 4 = -2 ( x , `y` ) This means, 4 is our new x-coordinate and -2 is our new y-coordinate. Therefore, your final answer is (4,-2). L (6, 2) → `L' (4,-2)`